A 50% alcohol solution is mixed with a 10% alcohol solution to create 8 ounces of a 32% alcohol solution. how much of each is needed.
thanks

Answer by lwsshak3(11628)

A 50% alcohol solution is mixed with a 10% alcohol solution to create 8 ounces of a 32% alcohol solution. how much of each is needed.
***
let x=amount of 50% solution to mix
8-x =amount of 10% solution to mix
50%x+10%(8-x)=32%*8
0.5x+0.8-0.1x=2.56
0.4x=1.76
x=4.4
8-x=3.6
amount of 50% solution needed=4.4 ounces
amount of 10% solution needed=3.6 ounces